OpenStudy (anonymous):

Solve the system of equations using the elimination method. x + 5y = -11 20y = -4x - 40

OpenStudy (anonymous):

answer please??

sam (.sam.):

Use this http://www.wolframalpha.com

OpenStudy (anonymous):

cool thanx but do u know the answer for this problem?

sam (.sam.):

x + 5y = -11 20y = -4x - 40 -------------------------- x + 5y = -11............................x4 4x+20y = - 40 -------------------------------- 4x+ 20y = -44 4x+20y = - 40

OpenStudy (anonymous):

?

OpenStudy (anonymous):

System is inconsistent. Rearrange the second to 4x+20y=-40, then divide by four. It will then say that negative ten is the same as negative eleven. This contradiction tells us the system is inconsistent, and has no solution.